OpenStudy (anonymous):

If one person can mow a lawn in 4 hours, and another person can mow the same lawn in 6 hours, how long will it take them to mow the lawn if they work together?

OpenStudy (amistre64):

the key is to figure out how much work each one does in an hour

OpenStudy (anonymous):

1/4 and 1/6

OpenStudy (amistre64):

very good, then 1/4 + 1/6 is mowed in one hour if they work together

OpenStudy (amistre64):

1/4 + 1/6 = 10/24 = 5/12 the neat thing about this is if we flip it over we get the time it takes :) it takes 12/5 hours to mow it together

OpenStudy (anonymous):

Oh, I remember this now! Thanks!

OpenStudy (amistre64):

youre welcome; its either that or you hire the 2 gents to mow the yard and record how long it took them ;)